# A Survey on Efficient Federated Learning Methods for Foundation Model Training

### v1

@@ 6 Related Work
While there are ample surveys that provide a broad perspective on FL Banabilah et al. (2022); Liu et al. (2022); Nguyen et al. (2021); Zhang et al. (2021), there are two closely related surveys to our work as they also focus on FMs.
@@ 7 Conclusions & Future Directions
In this paper, we survey the current landscape of computational and communication efficiency methods in FL systems and introduce a novel taxonomy based on the key techniques. While computational and communication efficiency have been separate topics in the past, they become closely intertwined as we start using FL systems to train FMs. Adjacent to our analysis of existing computational and communication efficiency methods in FL, we find the following three open questions that require attention from the FL community.
What are good and realistic evaluation strategies for generative downstream tasks in FL settings where we do not have control of data? Existing works focus on classification tasks. With the emergence of generative downstream tasks in FL systems, we require benchmarking tools that allow us to holistically validate the quality and efficiency of FL methods for FMs. It may be of particular interest to understand the notion of data heterogeneity in the context of generative tasks, i.e., how we define heterogeneous data splits.
How does hyperparameter optimization work for FMs in continuously evolving FL systems? While hyperparameter optimization in FL has been a key challenge, PEFT adds additional complexity. By the example of LoRA, we can systematically search for the optimal parameterization of $r$ while training in a centralized setting. In continuously evolving FL systems, parameterization needs to evolve over time. Furthermore, PEFT has been shown to be particularly sensitive to non-IID data, which further increases the complexity of hyperparameter optimization. SLoRA, for instance, shows that PEFT and FL can create good quality models, but to date, at significant costs in the form of long warm-up times.
How do PEFT and privacy in FL work together? Communication-efficient FL techniques, such as FedPM have been studied for their effect on privacy, but this is still an open topic for PEFT, PT, and IT. While it is proven that PEFT is more sensitive to data heterogeneity, the effects of perturbation through differential privacy are still subject to further studies. The same is true for PT and IT, as both techniques require precise prompts and instructions, respectively. As such, noise may have significantly negative effects here, as well.

### v3

@@ 6 Related Work
While there are ample surveys that provide a broad perspective on FL Li et al. (2023); Banabilah et al. (2022); Liu et al. (2022); Nguyen et al. (2021); Zhang et al. (2021); Aledhari et al. (2020), there are two closely related surveys to our work as they also focus on FMs.
@@ 7 Conclusions & Future Directions
In this paper, we survey the current landscape of computational and communication efficiency methods in FL systems and introduce a novel taxonomy based on the key techniques. While efficient FL methods have been separate topics on their own in the past, they become closely intertwined as we start using FL systems to leverage FMs. Consequently, the following three questions arise:
What are good and realistic evaluation strategies for generative downstream tasks in FL settings where we do not have control of data?
Fine-tuning generative FM requires high-quality data. However, we do not have access to data on the clients to monitor data quality before or during training. As such, estimating data quality is of utmost importance.
How does hyperparameter optimization work for FMs in continuously evolving FL systems?
While hyperparameter optimization in FL has been a key challenge, PEFT adds additional complexity. As such FL systems must adapt to the era of FMs by introducing adaptive parameterization for PEFT techniques that can account for changing environmental conditions.
We must develop an understanding of the interplay between PEFT and privacy in FL systems.
Communication-efficient FL techniques have been studied for their effect on privacy, but this is still an open topic for PEFT, PT, and IT. While it is proven that PEFT is more sensitive to data heterogeneity, the effects of perturbation through differential privacy are still subject to further studies. The same is true for PT and IT, as both techniques require precise prompts and instructions, respectively. As such, noise may have significantly negative effects here, as well.